OpenStudy (anonymous):

solve quadratic: x^2-4x-20=0

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-(-4)\pm \sqrt{-4^2-4(1)(-20)}}{2(1)}\]

OpenStudy (anonymous):

yeah. the answer's\[2+2\sqrt{6}\]

OpenStudy (anonymous):

the equation's wrong isn't it

OpenStudy (anonymous):

the quadratic equation I gave you. cause someone else solved that for me...

OpenStudy (anonymous):

yea i think the answer they gave you is wrong

OpenStudy (anonymous):

awesome. k, thanks though!

Directrix (directrix):

If x^2-4x-20=0 is the quadratic, then the roots are 2 - 2√6 and also 2 + 2√6.